OpenStudy (agentjamesbond007):

[Related Rates] Each side of a square is increasing at a rate of 6 cm/s. At what rate is the area of the square increasing when the area of the square is 16 cm^2?

OpenStudy (katrinakaif):

At a glance, this look accurate. But no guarantees; while since I have touched this.

OpenStudy (katrinakaif):

Agent, do you have an answer to this?

OpenStudy (ash2326):

Area= x^2 dx/dt=6 cm/s we have to find dA/dt when Area =16 or x=4 dA/dt=2x dx/dt dA/dt= 2*4*6 dA/dt=48 cm^2/sec area is increasing at the rate of 48cm^2/sec

OpenStudy (mimi_x3):

ooh..woops i thought the area was 64 lol